OpenStudy (anonymous):

What is the solution set of the following equation? -8x + x + 15 = -7x + 12 Ø {1/7} {all reals}

OpenStudy (anonymous):

combine like terms on each side of the equation, you get - 7x + 15 = - 7x + 12 Add 7x to both sides, you get 15 = 12 which is not a true statement, therefore, there is NO solution to the equation, or the empty set { } ...second choice.

OpenStudy (anonymous):

first choice, sorry

OpenStudy (anonymous):

-8x + x + 15 = -7x + 12 -7x+15=-7x+12 15=12(False) s.s= empty set